Four resistors having resistances of 20 Ω, 40 Ω, 60 Ω, and 80 Ω are connected in series across an ideal 50-V dc source. What is the current through each resistor?

a)0.75 A

b)2.0 A

c)4.0 A

d)0.25 A

e)0.50 A
Answer #1

Since resistors are connected in series,

R = R1 + R2 +R3 +R4

R = 20 + 40 + 60 +80 = 200 ohms

V = 50 V

V = IR

I = V/R = 50 /200 = 0.25 A

Since they are in series equal current passes through all the resistors

I = 0.25 A